Find three consecutive even integers such that the sum of the smallest number and twice the middle number is 20 more than the largest number.
Answer by Boreal(15235)

x, x+2, x+4 are the numbers
x+2(x+2)-20=x+4
3x-16=x+4
2x=20
x=10
10, 12, 14
10+24 is 20 more than 14.